When 35.0g of Ba(NO3)2 is reacted with excess Na2SO4 29.8g of BaSO4 is recovered by the chemist. What is the theoretical yield of BaSO4?
1
Expert's answer
2021-10-28T03:00:23-0400

Write down the reaction equation:

Ba(NO3)2 + Na2SO4 => BaSO4 + 2 NaNO3

From the equation it is clear that 1 mol of Ba(NO3)2 allows to obtain 1 mol of BaSO4

Molar mass of Ba(NO3)2 is 261.3 g/mol

Molar mass of BaSO4 is 233.4 g/mol

Find the theoretical yield of BaSO4:

"m"theor = "m_1*M_2\/M_1 = 35.0 * 233.4 \/ 261.3 = 31.3 (g)"

Answer: 31.3 g -- the theoretical yield of BaSO4.